OpenStudy (anonymous):

DIRECTIONS: solve each system by substitution. Problem: y=6x+16 3x-3y=-3 please help me.

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

help

OpenStudy (kirbykirby):

Since you know y in the first equation, substitute that into the "y" of the second equation: \(3x-3(6x+16)=-3\) Do you see that?

OpenStudy (anonymous):

yes

OpenStudy (kirbykirby):

So now you have just an equation with just on variable x, so you can solve for x: \(3x-18x-48=-3\) (I distributed the -3 to each term in the parentheses)

OpenStudy (anonymous):

ok i see that

OpenStudy (kirbykirby):

\(-15x-48=-3\) \(-15x=-3+48\) \(-15x=45\)\[ x=\frac{45}{-15}=-3\]

OpenStudy (anonymous):

let me write that down

OpenStudy (anonymous):

ok i got that written down

OpenStudy (kirbykirby):

Now you know the value of x... so you can plug in x = -3 into the first equation, and that will give you the value of y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how exatly to i do that?

OpenStudy (kirbykirby):

\(y=6x+16\) You know \(x=-3\), so just put it in where \(x\) is: \(y=6(-3)+16\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what do i do?

OpenStudy (kirbykirby):

just calculate \(6(-3)+16\) remeber that the notation \(6(-3)\) means \(6 \text{ times} -3\)

OpenStudy (anonymous):

ok thanks for taking your time to help me and have a great rest of the day

OpenStudy (kirbykirby):

yw same to you!

OpenStudy (anonymous):

ps you just gained a new fan

OpenStudy (kirbykirby):

:) awesome

OpenStudy (anonymous):

bye